OpenStudy (anonymous):

is the ordered pair a solution of the given system of equations ? (5,-5), 4x+9y=-25 3x-y=10

OpenStudy (anonymous):

I got 7=-25 and 9y=10 so it is a soltion of the system of equation

OpenStudy (anonymous):

what did you get it

OpenStudy (wolf1728):

Solving the equations A) 4x+9y=-25 B) 3x-y=10 Multiply B by 9 then add it to A A) 4x+9y=-25 B) 27x -9y = 90 31x = 65 x = 65/31 = 2.4193548387 x = 2.0967741935 The ordered pair states that x=5 which cannot be true and so the ordered pair is not a solution.

OpenStudy (anonymous):

I disagree with your answer (5,-5), 4x+9y=-2 and 3x-y=10 3x-y=10 3(3)-y(-y)=10 9+y=10 4x+9y=-25 9y=10 4(3)+(-5)=-25 12-5=-24 7=-25 So correct me where I did mistake

OpenStudy (wolf1728):

Solving for y 4*2.0967741935 + 9y = -25 9y = -25 - 8.3870967742 9y = -33.3870967742 y = -33.3870967742 / 9 y = -3.7096774194

OpenStudy (wolf1728):

shakur, Rather than plugging in the x and y values, I just solved for x and y. I didn't say whether you were wrong or right.

OpenStudy (anonymous):

Oh Okay I understand ,I just find it ,is not a solution